# spacing=nonfrench

### v1

@@ Abstract
Let $\Gamma$ be a bounded Jordan curve and $\Omega_{i},\Omega_{e}$ its two complementary components. For $1<p<\infty,\,s\in(0,1)$ we define $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ as the set of functions $f:\Gamma\to\mathbb{C}$ having harmonic extension $u$ respectively in $\Omega_{i}$ and $\Omega_{e}$ such that
@@
If $\Gamma$ is further assumed to be rectifiable we define $B_{p,p}^{s}(\Gamma)$ as the space of functions $f\in L^{p}(\Gamma)$ such that
@@
When $\Gamma$ is the unit circle these three spaces coincide with the homogeneous fractional Besov-Sobolev space. For a general rectifiable curve these spaces need not coincide and our first goal is to investigate the cases of equality: while the chord-arc property is the necessary and sufficient condition for equality in the classical case of $s=1/p,\,p\geq 2$, this is no longer the case for general $s\in(0,1)$. We show however that equality holds for radial-Lipschitz curves. In the general (possibly non-rectifiable) case we study boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and give conditions for equality of these trace-spaces that we then call $\mathcal{B}^{s}_{p,p}(\Gamma)$. Using Plemelj-Calderón property we further identify $\mathcal{B}^{s}_{p,p}(\Gamma)$ with the space of restrictions of a weighted Sobolev space of the plane. Finally we re-interpretate some of our results as the "almost"-Dirichlet principle in the spirit of Maz’ya.
Keywords: Cauchy integral, fractional Sobolev spaces, chord-arc curves, Lipschitz curves, Dirichlet spaces
@@ 1 Introduction
1.1 The Dirichlet Principle
In all this work we will be considering a bounded Jordan curve $\Gamma$ and call $\Omega_{i},\Omega_{e}$ respectively the interior and exterior connected components of its complement in the sphere. If $f$ is a continuous function on $\Gamma$ and $F$ a continuous extension of $f$ to $\overline{\Omega}_{i,e}$ which is furthermore assumed to be $C^{1}$ in $\Omega_{i,e}$ we call (2)(2-)energy the quantity
@@ 1.1 The Dirichlet Principle
The Dirichlet principle asserts that among all the extensions of $f$ the one with the lowest energy is the harmonic one. Notice that the (2)(2-)energy of the harmonic extension of a continuous function may be infinite, in which case every extension has infinite (2)(2-)energy. When $\Gamma=\mathbb{T}$, the unit circle, we define $\mathcal{D}(\mathbb{D}_{i})$, the Dirichlet space over the unit disk $\Omega_{i}=\mathbb{D}_{i}$ as the space of harmonic functions in $\mathbb{D}_{i}$ with finite (2)(2-)energy. This space coincides with the space of Poisson integrals of $L^{2}(\mathbb{T})$-functions such that
@@
The map $z\mapsto 1/\bar{z}$ is a reflection about the unit circle which is bi-Lipschitz from $\{1/2<|z|<2\}$ onto itself. It follows that if $u_{i}\in\mathcal{D}(\mathbb{D}_{i})$ is the Poisson integral of $f\in L^{2}(\mathbb{T})$ then the Poisson extension $u_{e}$ of $f$ in $\mathbb{D}_{e}=\overline{\mathbb{C}}\setminus\overline{\mathbb{D}}_{i}$ also has a finite energy. We may thus view the Dirichlet space as the subset of $L^{2}(\mathbb{T})$ of functions whose harmonic extension to $\overline{\mathbb{C}}\setminus\mathbb{T}$ has finite energy.
@@
It is not difficult to see that the above quantity, the square of the $H^{s}$-semi-norm, is equivalent (in particular, in the case of $s=1/2$ it is equal, up to a constant multiplicative factor,) to two other:
@@
The mention of Littlewood-Paley comes from the fact that for $s=0$ the Littlewood-Paley semi-norm is exactly the $L^{2}$-norm of $\mathsf{g}(f)$, the Littlewood-Paley function associated to $f$ [37].
@@
and that this space also coincides with $(I-\Delta)^{-s/2}(L^{2}(\mathbb{T}))$, namely the space of Bessel potentials of order $s$.
@@
so that, to avoid confusions, we will always use the Besov terminology at least for $p\neq 2$.
The main goal of this paper is to extend these notions to more general Jordan curves. The first restriction we want to impose on the Jordan curve $\Gamma$ is that we want it to share with $\mathbb{T}$ the reflection property, that is we want the existence of a reflection across $\Gamma$ which is bi-Lipschitz in a neighborhood of $\Gamma$. The curves satisfying this property are well-known: they are the quasicircles, i.e., the quasiconformal images of circles [1]. Geometrically, the Jordan curve $\Gamma$ is a quasicircle if and only if there exists $C>1$ such that
@@
A quasicircle need not be smooth, and the Hausdorff dimension of a quasicircle can take any value in $[1,2)$. As an example, the Von Koch snowflake curve is a quasicircle without any tangent and whose Hausdorff dimension, which coincides with its Minkowski dimension, is $\log 4/\log 3$.
@@
1.2 Statement of the Main Theorems
Let $\Gamma$ be a quasicircle and $\Omega_{i,e}$ the inner and outer connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. Suppose $\Omega$ is one of these two components and correspondingly $\mathbb{D}$ is one of $\mathbb{D}_{i}$ and $\mathbb{D}_{e}$. We define
$\mathcal{B}_{p,p}^{s}(\Omega)=\{u\;\mathrm{harmonic\;in}\;\Omega:\;\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}:=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\}.$
We define for functions in these spaces and for some values of p,sp,\,s their boundary values on $\Gamma$, which will allow us to identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with a space of functions defined on $\Gamma$ that we denote by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$.
In order to define this space we distinguish three cases, namely, s>1/p,s=1/p,s<1/ps>1/p,\,s=1/p,\,s<1/p.
@@ 1.2 Statement of the Main Theorems
    Since this latter space is included in the harmonic Hardy space $h^{p}(\mathbb{D})$ we conclude that $f$ has a well-defined boundary value in $L^{p}(\omega,\Gamma)$, where $\omega$ is the harmonic measure for $\Omega$.
@@
    Case $s<1/p$: we proceed here differently. Let $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\;f\in C_{c}^{\infty}(\mathbb{C})\}$, $C_{c}^{\infty}(\mathbb{C})$-functions being $C^{\infty}$-functions with compact support in $\mathbb{C}$. Let $\mathcal{D}_{p}^{s}(\Gamma)$ be the set of functions in $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ such that their harmonic extensions to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and we denote the set of such extensions by $\mathcal{D}_{p}^{s}(\Omega_{i,e})$. We define the space of boundary values $\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\!\to\!\Gamma)$ as the abstract closure of this space, i.e., equivalence classes of Cauchy sequences in $\mathcal{D}_{p}^{s}(\Gamma)$, equipped with the $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$-norm. Notice that this procedure only defines boundary values for the closures of $\mathcal{D}_{p}^{s}(\Gamma)$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, which may not be the full space. However, the space $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ in this case is quite problematic, and the closure of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is actually a more robust variant of this space. .
For all these cases we define
@@
equipped with the natural norm. One of our main tasks will be to examine when these three spaces coincide. The case 2 above is particularly interesting especially since the two harmonic measures may be mutually singular. This drawback has been ruled by [36] and [5] by exhibiting a change of variable realizing a "transmission" operator between the spaces on both sides which happens to be an isomorphism, thus showing the equality of the three spaces for all quasicircles.
When $\Gamma$ is furthermore assumed to be rectifiable we define, following Douglas [14],
$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\;\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}:=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
Here and in what follows, $|dz|$ denotes the arc-length measure. If $\Gamma$ is the unit circle the three spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma),\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),B_{p,p}^{s}(\Gamma)$ coincide (see p.151 in [38]), but they have no reason to coincide in general. Finding cases of equality is one of our goals. In order to state the first result in this direction, we need some definitions:
@@
where $\gamma_{j},\,j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$. Notice that chord-arc curves are in particular rectifiable quasicircles (but the converse is not true).
We will also need the notion of radial-Lipschitz curves. These are the curves given in polar coordinates by the equation $z(\theta)=r(\theta)e^{i\theta}$ where $r:\mathbb{R}\to(0,+\infty)$ is $2\pi$-periodic and Lipschitz continuous. The radial-Lipschitz curves are in particular chord-arc.
In [44] and [45] the present authors have proven
@@
Let $p>1$. If $\Gamma$ is chord-arc then
@@
The methods we used for the proof of this theorem depend heavily on the special features of the case $s=1/p$, namely, conformal invariance and the existence of a transmission operator.
In this paper our main results in this direction are Theorems 5.5, 5.7, 5.8 combined with the claims in appendix that loosely speaking say that
Theorem 1.2.
If $\Gamma$ is a radial-Lipschitz curve and $p=2$, $s\in(0,1)$ then
@@
This can be generalized to any $p\in(1,\infty)$ under a slightly different assumption related to $p$.
We will also prove a result in this regard for general quasicircles. Before stating it we need to introduce some notions:
Astala [3] has introduced the notion of "Minkowski content" $h(\Gamma)$: the proper definition will be recalled below but let us just mention here that $h(\Gamma)$ is the Minkowski dimension if $\Gamma$ is a self-similar fractal set (as is the Koch snowflake curve for example).
The Muckenhoupt weights $A_{p}(\mathbb{C}),\,p>1$, are the weights $\omega$ such that the Hardy-Littlewood maximal function is bounded on $L^{p}(\omega,\mathbb{C})$. Let $\omega$ be such a weight. We define the weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ as being the space of tempered distributions $f$ such that $f$ is locally integrable and, in the sense of distributions, $\nabla f\in L^{p}(\omega,\mathbb{C})$.
@@
If $\Gamma$ is a quasicircle and
$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
where $K_{*}\leq 2$ is a geometric quantity depending only on $\Gamma$ to be defined in section $3$ then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ (1.3)
where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$.
Theorem 1.3 allows us to improve theorem 1.2 for certain values of $p,s$:
Corollary 1.4.
If $\Gamma$ is chord-arc and $p>2,s\in[1-\frac{1}{p},\frac{p+K_{*}-1}{K_{*}p})$ then (1.3) holds.
Another corollary is an "almost-Dirichlet principle" generalizing, in dimension $2$ a result of Maz’ya simplified by Mironescu-Russ ([29], [30]). We say that the quasidisk $\Omega$ satisfies the almost-Dirichlet principle for $p,s$ if any continuous function $f:\Gamma\to\mathbb{C}$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$ and satisfies
@@
Corollary 1.5.
Under the assumption of Theorem 1.3 the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$.
One of the tools for proving Theorem 1.3 will be the use of Plemelj-Calderón property which is of independent interest.
1.3 Plemelj-Calderón Problem
It is an old problem, given a bounded Jordan curve $\Gamma$ in the plane and a (complex valued) function $f$ defined on $\Gamma$, to find two holomorphic functions $G_{i},G_{e}$ defined on the interior-connected and the exterior-connected components of $\Gamma$ such that
$f=G_{i}|_{\Gamma}+G_{e}|_{\Gamma},$
the boundary traces being defined in some sense.

### v2

@@ Abstract
Let $\Gamma$ be a bounded Jordan curve and $\Omega_{i},\Omega_{e}$ its two complementary components. For $p\in(1,\infty),\,s\in(0,1)$ we define the two spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ as the set of harmonic functions $u$ respectively in $\Omega_{i}$ and $\Omega_{e}$ such that
@@
When it is possible to identify these spaces with spaces of functions on the boundary (trace spaces), we address the question of their equality. When $\Gamma$ is the unit circle, these two spaces coincide with homogeneous fractional Besov-Sobolev spaces and the framework of quasicircles appears to be an appropriate generalization. In this framework, we study the boundedness of the Plemelj-Calderón operator and apply the results to show that for some values of $p,s$, if the two spaces coincide, they are restrictions to $\Gamma$ of some weighted Sobolev space.
If $\Gamma$ is further assumed to be rectifiable, we define $B_{p,p}^{s}(\Gamma)$ as the space of functions $f\in L^{p}(\Gamma)$ such that
@@
Again, these spaces coincide with the homogeneous fractional Besov-Sobolev spaces for the unit circle. While the chord-arc property is the necessary and sufficient condition for the equality
$\mathcal{B}_{p,p}^{s}(\Omega_{i})=\mathcal{B}_{p,p}^{s}(\Omega_{e})=B_{p,p}^{s}(\Gamma)$
in the case of $s=1/p,\,p\geq 2$, this is no longer the case for general $s\in(0,1)$. However, we show that equality holds for radial-Lipschitz curves. Finally, we re-interpretate some of our results as some "almost"-Dirichlet principle in the spirit of Maz’ya.
Keywords: Cauchy integral, fractional Sobolev spaces, Besov spaces, quasicircles, chord-arc curves, Lipschitz curves, Dirichlet spaces
@@ 1 Introduction
1.1 Dirichlet principle
In all this work, we will consider a bounded Jordan curve $\Gamma$ and call $\Omega_{i},\Omega_{e}$ the interior and exterior connected components of its complement in the sphere. If $f$ is a continuous function on $\Gamma$ and $F$ a continuous extension of $f$ to $\overline{\Omega}_{i,e}$, which is furthermore assumed to be $C^{1}$ in $\Omega_{i,e}$, we call (2)(2-)energy the quantity
@@ 1.1 Dirichlet principle
The Dirichlet principle states that among all the extensions of $f$ the one with the lowest energy is the harmonic one. Notice that the (2)(2-)energy of the harmonic extension of a continuous function may be infinite, in which case every extension has infinite (2)(2-)energy. When $\Gamma=\mathbb{T}$, the unit circle, we define $\mathcal{D}(\mathbb{D}_{i})$, the Dirichlet space on the unit disk $\Omega_{i}=\mathbb{D}_{i}$ as the space of harmonic functions in $\mathbb{D}_{i}$ with finite (2)(2-)energy. This space coincides with the space of Poisson integrals of $L^{2}(\mathbb{T})$-functions such that
@@
The map $z\mapsto 1/\bar{z}$ is a reflection about the unit circle that is bi-Lipschitz from $\{1/2<|z|<2\}$ onto itself. It follows that if $u_{i}\in\mathcal{D}(\mathbb{D}_{i})$ is the Poisson integral of $f\in L^{2}(\mathbb{T})$ then the Poisson extension $u_{e}$ of $f$ in $\mathbb{D}_{e}=\overline{\mathbb{C}}\setminus\overline{\mathbb{D}}_{i}$ also has a finite energy.
@@
It is not difficult to see that the above quantity, the square of the $H^{s}$-semi-norm, is equivalent (in particular, in the case of $s=1/2$ it is equal, up to a constant multiplicative factor) to two other:
@@
The mention of Littlewood-Paley comes from the fact that for $s=0$ the Littlewood-Paley semi-norm is exactly the $L^{2}$-norm of $\mathsf{g}(f)$, the Littlewood-Paley function associated with $f$ (see [44]).
@@
and this space also coincides with $(I-\Delta)^{-s/2}(L^{2}(\mathbb{T}))$, that is, the space of Bessel potentials of order $s$.
@@
so that, to avoid confusion, we will use the Besov terminology. The spaces $B_{p,p}^{s}(\mathbb{T})$, $s\in(0,1)$, $p\in(1,\infty)$, are "intermediate" spaces between two Sobolev spaces, so they are also called fractional Sobolev spaces in the literature. Throughout this paper, we call them (along with their generalizations to the more general Jordan curves) fractional Besov-Sobolev spaces.
The main goal of this paper is to extend these notions to more general Jordan curves. The first restriction on the Jordan curve $\Gamma$ is that we want it to share with $\mathbb{T}$ the reflection property, that is, the existence of a reflection across $\Gamma$ which is bi-Lipschitz in a neighborhood of $\Gamma$. The curves satisfying this property are well-known: they are the quasicircles, i.e., the quasiconformal images of circles [1]. Geometrically, the Jordan curve $\Gamma$ is a quasicircle if and only if there exists $C>1$ such that
@@
A quasicircle does not necessarily need to be smooth, and the Hausdorff dimension of a quasicircle can take any value in $[1,2)$. As an example, the Von Koch snowflake curve is a quasicircle without any tangent, and its Hausdorff dimension, which coincides with its Minkowski dimension, is $\log 4/\log 3$.
@@
1.2 Statement of main theorems
Let $\Gamma$ be a quasicircle and $\Omega_{i,e}$ the inner and outer connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. Suppose $\Omega$ is one of these two components and, correspondingly, $\mathbb{D}$ is one of $\mathbb{D}_{i}$ and $\mathbb{D}_{e}$. We define, for $p>1$ and $0<s<1$,
$\mathcal{B}_{p,p}^{s}(\Omega)=\{u\;\mathrm{harmonic\;in}\;\Omega:\;\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\}.$
We now define for functions in these spaces and for some values of p,sp,\,s their boundary values on $\Gamma$, which will allow us to identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with a space of functions defined on $\Gamma$ that we denote by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$.
In order to define this space, we distinguish three cases, namely s>1/p,s=1/p,s<1/ps>1/p,\,s=1/p,\,s<1/p.
@@ 1.2 Statement of main theorems
    Since the latter space is included in the harmonic Hardy space $h^{p}(\mathbb{D})$, we conclude that $f$ has a well-defined boundary value in $L^{p}(\omega,\Gamma)$, where $\omega$ is the harmonic measure for $\Omega$.
@@
    Case $s<1/p$: we proceed here differently. Let $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\;f\in C_{c}^{\infty}(\mathbb{C})\}$, $C_{c}^{\infty}(\mathbb{C})$-functions being $C^{\infty}$-functions with compact support in $\mathbb{C}$. Let $\mathcal{D}_{p}^{s}(\Gamma)$ be the set of functions in $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ such that their harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and we denote the set of such extensions by $\mathcal{D}_{p}^{s}(\Omega_{i,e})$. We define the space of boundary values $\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\!\to\!\Gamma)$ as the abstract closure of this space, i.e., the equivalence classes of Cauchy sequences in $\mathcal{D}_{p}^{s}(\Gamma)$, equipped with the $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$-norm. Notice that this procedure only defines boundary values for the closures of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, which may not be the full space. However, the space $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ in this case is quite problematic, and the closure of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is actually a more robust variant of this space.
For all these cases, we define
@@
equipped with the natural norm. The case 2 above is particularly interesting, especially since the two harmonic measures may be mutually singular. This drawback has been ruled out by [40] and [5] by exhibiting a change of variables realizing a "transmission" operator between the spaces on both sides which happens to be an isomorphism, thus showing the equality of the three spaces for all quasicircles.
When $\Gamma$ is furthermore assumed to be rectifiable, we define, following Douglas [14],
$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\;\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
Here and in what follows, $|dz|$ denotes the arc-length measure. If $\Gamma$ is the unit circle, the three spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma),\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),B_{p,p}^{s}(\Gamma)$ coincide (see p.151 in [45]), but they have no reason to coincide in general. Finding cases of equality is one of our goals. In order to state the first result in this direction, we need some definitions:
@@
where $\gamma_{j},\,j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$. Notice that chord-arc curves are, in particular, rectifiable quasicircles (but the converse is not true).
We will also need the notion of radial-Lipschitz curves. These are the curves given in polar coordinates by the equation $z(\theta)=r(\theta)e^{i\theta}$ where $r:\mathbb{R}\to(0,+\infty)$ is $2\pi$-periodic and Lipschitz continuous (The norm of Lipschitz curve will be defined in Section 5.). The radial-Lipschitz curves are in particular chord-arc.
@@
Let $p>1$. If $\Gamma$ is a chord-arc, then
@@
The second statement was first proved in [50] and [51] by the present authors. For the proof of the first statement, we refer to [27, 29, 28] (see also [41, 40, 43]). The special features of the case $s=1/p$ lie in (and the proof of this theorem depends primarily on) conformal invariance of spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ and the existence of a transmission operator.
In Section 5 we will prove our main results in this direction (see Theorems 5.5 and 5.7) that loosely speaking say that
Theorem 1.2 (see Corollary 5.10).
If $\Gamma$ is a radial-Lipschitz curve with norm $M$ then
@@
for any $s\in(0,1)$ and $p\in\left(1+\frac{2\arctan M}{\pi},1+\frac{\pi}{2\arctan M}\right)$. In particular, this interval converges to $(1,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
In Section 3, we will also prove a result in this regard for general quasicircles. Before stating it, we need to introduce some notions.
Astala [3] has introduced the notion of "Minkowski content" $h(\Gamma)$: the proper definition will be recalled below but let us just mention here that $h(\Gamma)$ is the Minkowski dimension if $\Gamma$ is a self-similar fractal set (as is the Von Koch snowflake curve, for example), and that chord-arc curves are exactly quasicircles with $h(\Gamma)=1$.
The Muckenhoupt weights $A_{p},\,p>1$, in $\mathbb{C}$ are the weights $\omega$ such that the Hardy-Littlewood maximal function is bounded in $L^{p}(\omega,\mathbb{C})$. Let $\omega$ be such a weight. We define the weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ as the space of tempered distributions $f$ such that $f$ is locally integrable and, in the sense of distributions, $\nabla f\in L^{p}(\omega,\mathbb{C})$.
@@
Let $\Gamma$ be a quasicircle and $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$. If
$p>2,\quad\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}<s<1-\frac{h(\Gamma)-1}{p}$ (1.3)
(the latter reduces to $1/2<s<1$ when $\Gamma$ is chord-arc) and
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$ (1.4)
then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}.$ (1.5)
We remark here that in this theorem the condition (1.3) is used to guarantee that $W^{1,p}(\omega,\mathbb{C})$ consists of continuous functions so that its restriction on $\Gamma$ makes sense, and the condition (1.3) can also ensure that $\omega$ is an $A_{p}$-weight.
Examples of pairs $(p,s)$ such that equation (1.4) holds are $(p,1/p)$ and $(\infty,s)$ with $s<1/2$. If one could interpolate between these special spaces, one could obtain much more spaces satisfying (1.4). We will investigate this possibility in our forthcoming paper. Another class of quasicircles satisfying (1.4) is given by Theorem 1.2: it follows that in this case
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=B_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}.$
A corollary of this study is an "almost-Dirichlet principle" generalizing, in dimension $2$ a result of Maz’ya simplified by Mironescu-Russ ([32], [33]). We say that the quasidisk $\Omega$ satisfies the almost-Dirichlet principle for $p,s$ if any continuous function $f:\Gamma\to\mathbb{C}$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$ and satisfies
@@
Corollary 1.4 (see Corollary 5.12).
If $\Gamma$ is a radial-Lipschitz curve with norm $M$, then $\mathcal{B}_{p,p}^{s}(\Omega)$ has the almost-Dirichlet property if
$\frac{1}{2}<s<1\quad\text{and}\quad 2<p<1+\frac{\pi}{2\arctan M},$
that converges to $(2,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
One of the tools for proving Theorem 1.3 will be Plemelj-Calderón property, which is of independent interest.
1.3 Plemelj-Calderón problem: statement of two more main theorems
It is an old problem, given a bounded Jordan curve $\Gamma$ in the plane and a (complex valued) function $f$ defined on $\Gamma$, to find two holomorphic functions $G_{i},G_{e}$ defined in the interior and the exterior-connected components $\Omega_{i,e}$ of $\Gamma$ such that
$f=G_{i}|_{\Gamma}+G_{e}|_{\Gamma}.$
Here, the boundary traces are defined in some sense.